OpenStudy (anonymous):

Solve the logarithm equations: log(1 - X) + log x = log(x-3)

OpenStudy (raden):

first, log a + log b = log (ab). so log (1-x) + log x = log (1-x)x now we have equation : log (x-1)x = log (x-3) take (x-1)x = x-3 then solve for x